# Minimal complete tri-hypercharge theories of flavour

### v1

@@ 1 Introduction
Table 1: Field content of the low-energy (EFT) tri-hypercharge theory. $q_{i}$ and $\ell_{i}$ (where $i=1,2,3$) are left-handed (LH) $SU(2)_{L}$ doublets of chiral quarks and leptons, while $u_{i}^{c}$, $d_{i}^{c}$ and $e_{i}^{c}$ are the CP conjugated right-handed (RH) quarks and leptons (so that they become left-handed). Two singlet neutrinos $\nu_{1,2}^{c}$, also taken as left-handed fields, will be relevant for the origin of neutrino masses. $H_{3}^{u,d}$ are $SU(2)_{L}$ Higgs doublets, while $\phi_{q12,\ell 12,q23,\ell 23}$ are the so-called scalar hyperons which spontaneously break the three gauge hypercharges down to SM hypercharge (the diagonal subgroup $Y=Y_{1}+Y_{2}+Y_{3}$).
@@ 2 Minimal Models
$+\begin{pmatrix}\ell_{1}&\ell_{2}&\ell_{3}\end{pmatrix}\begin{pmatrix}{\displaystyle\frac{\tilde{\phi}_{\ell 12}}{\Lambda_{1}}\frac{\tilde{\phi}_{\ell 23}}{\Lambda_{2}}}&{\displaystyle\frac{\phi_{\ell 12}}{\Lambda_{1}}\frac{\tilde{\phi}_{\ell 23}}{\Lambda_{2}}}&{\displaystyle\frac{\phi_{\ell 12}}{\Lambda_{1}}\frac{\phi_{\ell 23}}{\Lambda_{2}}}\\ \vspace{-0.14cm}&\vspace{-0.14cm}&\vspace{-0.14cm}\\ {\displaystyle\frac{\tilde{\phi}_{\ell 12}}{\Lambda_{1}}\frac{\tilde{\phi}_{\ell 12}}{\Lambda_{1}}\frac{\tilde{\phi}_{\ell 23}}{\Lambda_{2}}}&{\displaystyle\frac{\tilde{\phi}_{\ell 23}}{\Lambda_{2}}}&{\displaystyle\frac{\phi_{\ell 23}}{\Lambda_{2}}}\\ \vspace{-0.14cm}&\vspace{-0.14cm}&\vspace{-0.14cm}\\ {\displaystyle\frac{\tilde{\phi}_{\ell 12}^{2}}{\Lambda_{1}}\frac{\tilde{\phi}_{\ell 23}^{2}}{\Lambda_{2}}}&{\displaystyle\frac{\tilde{\phi}_{\ell 23}^{2}}{\Lambda_{2}^{2}}}&1\end{pmatrix}\begin{pmatrix}e_{1}^{c}\\ e_{2}^{c}\\ e_{3}^{c}\end{pmatrix}H_{3}^{d}\,+\mathrm{h.c.}$ (15)
@@ 2.1 Model 1
Table 2: Scalar and vector-like fermion content of Model 1. Notice that for vector-like fermions (highlighted in yellow), only the left-handed fields $U,D,E,N$, with the same electric charges as the SM fermions $u^{c},d^{c},e^{c},\nu^{c}$ in Table 1, are displayed, while their vector-like partners with conjugate quantum numbers are implicitly assumed to be present but not explicitly shown.
@@
provides an excellent explanation of the SM flavour structure, which is translated to three simple and physical scales of new physics (NP) above electroweak symmetry breaking, as shown in Fig. 2. Indeed, we can explain the whole flavour structure of the SM in terms of just three small parameters,
@@
where, thanks to the non-generic relation of Eq. (25), testing experimentally any of these scales effectively tests all of them (and in particular, setting bounds on any of them also sets bounds on the rest). Notice that our model then translates the complicated flavour structure of the SM into just three simple and physical scales of new physics above electroweak symmetry breaking.
@@ 2.2 Model 2
${\displaystyle\frac{\langle{\phi}_{\ell 23}\rangle}{M_{U_{23}}}}\sim\lambda^{3}\,,\qquad{\displaystyle\frac{\langle{\phi}_{q23}\rangle}{M_{U_{23}}}}\sim\lambda^{2}\,,\qquad\frac{\langle{\phi}_{\ell 12}\rangle}{M_{U_{12,13}}}\sim\lambda^{2}\,,\qquad\frac{\langle{\phi}_{q12}\rangle}{M_{U_{12,13}}}\sim\lambda\,.$ (38)
@@ 3 Neutrinos
The above result is similar to what one would obtain in a minimal type-I seesaw extension of the SM. Indeed the texture above predicts two very light active neutrinos (and one massless) if $M_{11}\approx 10^{15}$ GeV, and is capable to explain large neutrino mixing by fitting the $\mathcal{O}(1)$ coefficients to existing data. Notably, this implementation of the seesaw mechanism requires no addition of extra scalars beyond those already introduced to explain the flavour structure of the charged fermion sector. Moreover, there is no need to impose small couplings nor assuming that the scales $v_{12}$ or $v_{23}$ are very heavy in order to achieve small neutrino masses (meaning that such scales can be in the multi-TeV region), as is commonly assumed in the literature.
@@ 4.1 Gauge boson couplings and complete fermion mixing
$g_{12}=\frac{g_{1}g_{2}}{\sqrt{g_{1}^{2}+g_{2}^{2}}},\qquad g_{Y}=\frac{g_{12}g_{3}}{\sqrt{g_{12}^{2}+g_{3}^{2}}}\,.$ (59)
@@ 4.2 Bounds from FCNCs
In contrast, the $Z^{\prime}_{23}$ boson is mostly protected from the most dangerous FCNCs by an approximate $U(2)^{5}$ flavour symmetry. The exception to this appears in Model 1 only, where we find that $Z^{\prime}_{23}$ mediates a 1-loop contribution to $\mu\rightarrow e\gamma$ with $E_{23}$ running in the loop, as shown in Fig. 5 (right), that appears due to the mixing of the $E_{23}$ VL fermion [^6] This is a consequence of non-minimal breaking of the approximate $U(2)^{5}$ flavour symmetry by the messengers $F_{23}$ in Model 1. This is avoided in the (less minimal) model of FernandezNavarro:2023hrf , where the leading breaking of $U(2)^{5}$ is done minimally, SM-like via the VL fermion $Q_{3}\sim(\mathbf{3,2})_{(0,0,\frac{1}{6})}$ and heavy Higgs doublets $H_{2}^{u,d}\sim(\mathbf{1,2})_{(0,\frac{1}{2},0)}$. Alternatively, in Model 2 the $U(2)^{5}$ symmetry is broken minimally in the down and charged lepton sectors, and non-minimally in the up sector where experimental bounds are much weaker. with $\mu$ and $e$. This contribution to $\mu\rightarrow e\gamma$ is chirally enhanced by a factor $M_{E_{23}}/m_{\mu}\sim 10^{7}$ but also suppressed by small mixing angles $s_{26}^{e_{R}}\sim\mathcal{O}(\langle\phi_{\ell 23}\rangle/M_{E_{23}}\sim\lambda^{3})$ and $s_{16}^{e_{L}}\sim\mathcal{O}(s_{12}^{e_{L}}\langle H_{3}^{d}\rangle/M_{E_{23}}\sim 10^{-7})$. As shown in Table 6 and Fig. 6(a), this contribution leads to a bound in the few TeV region, stronger than the bounds obtained from $3\rightarrow 2$ flavour transitions such as $B_{s}$ meson mixing or $B_{s}\rightarrow\mu\mu$. In Fig. 6(a) we also depict in green the region of parameter space that could explain the anomaly in $b\rightarrow s\nu\nu$ transitions recently observed by Belle II Belle-II:2023esi , although this region is excluded by electroweak precision and by the leading flavour-violating observables. Similarly, we expect significant enhancements of $b\rightarrow s\mu\mu$ or $b\rightarrow s\tau\tau$ (see e.g. Capdevila:2017iqn ; Alguero:2023jeh ; Bordone:2023ybl ) to be in conflict with electroweak precision and with the leading flavour observables.
@@
This bound is well compatible with the model-independent phenomenology of $Z^{\prime}_{23}$, which is dominated by the production of dilepton tails at the LHC and the modification of electroweak precision observables via mixing [^7] As shown in FernandezNavarro:2023rhv , this mass mixing is present even if kinetic mixing among the three gauge $U(1)$ factors is negligible, as it happens when they originate from a semi-simple embedding like in FernandezNavarro:2023hrf . with the SM $Z$ boson FernandezNavarro:2023rhv ; Davighi:2023evx . In particular, assuming all Yukawa couplings of the theory to be of $\mathcal{O}(1)$, and assuming $g_{3}(10\,\mathrm{TeV})\simeq 0.6$ as suggested by a GUT origin FernandezNavarro:2023hrf , we obtain the same bound $M_{Z^{\prime}_{23}}>5\,\mathrm{TeV}$ from both $\mu\rightarrow e\gamma$ and from dilepton searches at the LHC. Therefore, to be conservative we set $v_{23}\sim\mathcal{O}(10\,\mathrm{TeV})$ and, thanks to the predictive relation in Eq. (25), this fixes the the whole spectrum of both models as in Fig. 2, and in particular the mass of $Z^{\prime}_{12}$ as $M_{Z^{\prime}_{12}}\sim v_{12}\sim\mathcal{O}(10^{3}\,\mathrm{TeV})$, well compatible with all the bounds shown in Table 6.
@@ Appendix C EFT formalism for FCNCs
Now we apply a Fierz transformation to the LR operators in order to match to the basis of Refs. UTfit:2007eik ; Bona:2024bue ; UTfit2023 , from which we take the bounds, obtaining
$\mathcal{L}_{\mathrm{EFT}}\supset-\left[C_{1}^{sd}(\bar{s}_{L}\gamma_{\mu}d_{L})^{2}-2C_{LR}^{sd}(\bar{s}_{L}^{\alpha}d_{R}^{\beta})(\bar{s}_{R}^{\beta}d_{L}^{\alpha})+C_{1}^{cu}(\bar{c}_{L}\gamma_{\mu}u_{L})^{2}-2C_{LR}^{cu}(\bar{c}_{L}^{\alpha}u_{R}^{\beta})(\bar{c}_{R}^{\beta}u_{L}^{\alpha})\right]\,.$ (85)

### v2

@@ 1 Introduction
Table 1: Field content of the low-energy (EFT) tri-hypercharge theory. $q_{i}$ and $\ell_{i}$ (where $i=1,2,3$) are left-handed (LH) $SU(2)_{L}$ doublets of chiral quarks and leptons, while $u_{i}^{c}$, $d_{i}^{c}$ and $e_{i}^{c}$ are the CP conjugate [^1] For a given fermion $\Psi$ that consists of two 4-component chiral spinors $\Psi_{L}=(\psi_{L},0)^{\mathrm{T}}$ and $\Psi_{R}=(0,\psi_{R})^{\mathrm{T}}$, where $\psi_{L}$ and $\psi_{R}$ are 2-component (Weyl) spinors, one may get rid of chiral indices by defining $\psi\equiv\psi_{L}$ and by taking the CP conjugate of $\psi_{R}$, i.e. $\psi^{c}\equiv CP\psi_{R}(CP)^{-1}$, which is by construction a left-handed field. Then one can write the theory in terms of $\psi$ and $\psi^{c}$ only. For example, fermion bilinears among both notations are related by $\overline{\Psi}_{R}\Psi_{L}\rightarrow\psi^{c}\psi$. We also refer the reader to Dreiner:2008tw ; Martin:2012us for extensive reviews on 2-component spinor notation and techniques.right-handed (RH) quarks and leptons (so that they become left-handed). Two singlet neutrinos $\nu_{1,2}^{c}$, also taken as left-handed fields, will be relevant for the origin of neutrino masses. $H_{3}^{u,d}$ are $SU(2)_{L}$ Higgs doublets, while $\phi_{q12,\ell 12,q23,\ell 23}$ are the so-called scalar hyperons which spontaneously break the three gauge hypercharges down to SM hypercharge (the diagonal subgroup $Y=Y_{1}+Y_{2}+Y_{3}$).
@@ 2 Minimal Models
$+\begin{pmatrix}\ell_{1}&\ell_{2}&\ell_{3}\end{pmatrix}\begin{pmatrix}{\displaystyle\frac{\tilde{\phi}_{\ell 12}}{\Lambda_{1}}\frac{\tilde{\phi}_{\ell 23}}{\Lambda_{2}}}&{\displaystyle\frac{\phi_{\ell 12}}{\Lambda_{1}}\frac{\tilde{\phi}_{\ell 23}}{\Lambda_{2}}}&{\displaystyle\frac{\phi_{\ell 12}}{\Lambda_{1}}\frac{\phi_{\ell 23}}{\Lambda_{2}}}\\ \vspace{-0.14cm}&\vspace{-0.14cm}&\vspace{-0.14cm}\\ {\displaystyle\frac{\tilde{\phi}_{\ell 12}}{\Lambda_{1}}\frac{\tilde{\phi}_{\ell 12}}{\Lambda_{1}}\frac{\tilde{\phi}_{\ell 23}}{\Lambda_{2}}}&{\displaystyle\frac{\tilde{\phi}_{\ell 23}}{\Lambda_{2}}}&{\displaystyle\frac{\phi_{\ell 23}}{\Lambda_{2}}}\\ \vspace{-0.14cm}&\vspace{-0.14cm}&\vspace{-0.14cm}\\ {\displaystyle\frac{\tilde{\phi}_{\ell 12}^{2}}{\Lambda^{2}_{1}}\frac{\tilde{\phi}_{\ell 23}^{2}}{\Lambda^{2}_{2}}}&{\displaystyle\frac{\tilde{\phi}_{\ell 23}^{2}}{\Lambda_{2}^{2}}}&1\end{pmatrix}\begin{pmatrix}e_{1}^{c}\\ e_{2}^{c}\\ e_{3}^{c}\end{pmatrix}H_{3}^{d}\,+\mathrm{h.c.}$ (15)
@@ 2.1 Model 1
Table 2: Scalar and vector-like fermion content of Model 1. Notice that for vector-like fermions (highlighted in yellow), only the left-handed fields $U,D,E,N$, with the same electric charges as the chiral fermions $u^{c},d^{c},e^{c},\nu^{c}$ in Table 1, are displayed, while their vector-like partners with conjugate quantum numbers are implicitly assumed to be present but not explicitly shown.
@@
provides an excellent explanation of the SM flavour structure, which is translated to three simple and physical scales [^4] In principle, the masses of VL fermions are arbitrary parameters and finding a dynamical mechanism that generates them at the scales of Fig. 2 is not straightforward. The VL fermions could be made chiral under an extended symmetry, in such a way that the VL fermions get their mass when this symmetry is broken, similar to how right-handed neutrinos get mass in Majoron models where lepton number is spontaneously broken. However, the implementation of such a mechanism in our theory lies beyond the scope of this work. of new physics (NP) above electroweak symmetry breaking, as shown in Fig. 2. Indeed, we can explain the whole flavour structure of the SM in terms of just three small parameters,
@@
where, thanks to the non-generic relation of Eq. (25), testing experimentally any of these scales effectively tests all of them (and in particular, setting bounds on any of them also sets bounds on the rest). Notice that our model then translates the complicated flavour structure of the SM into just three simple and physical scales of new physics above electroweak symmetry breaking, as shown in Fig. 2.
@@ 2.2 Model 2
${\displaystyle\frac{\langle{\phi}_{\ell 23}\rangle}{M_{U_{23}}}}\sim\lambda^{3}\,,\qquad{\displaystyle\frac{\langle{\phi}_{q23}\rangle}{M_{U_{23}}}}\sim\lambda^{2}\,,\qquad\frac{\langle{\phi}_{\ell 12}\rangle}{M_{U_{12,13}}}\sim\lambda^{2}\,,\qquad\frac{\langle{\phi}_{q12}\rangle}{M_{U_{12,13}}}\sim\lambda\,,$ (38)
@@ 3 Neutrinos
This result is similar to what one would obtain in a minimal type-I seesaw extension of the SM. Indeed the texture above predicts two very light active neutrinos (and one massless) if $M_{11}\approx 10^{15}$ GeV, and is capable to explain large neutrino mixing by fitting the $\mathcal{O}(1)$ coefficients to existing data. Notably, this implementation of the seesaw mechanism requires no addition of extra scalars beyond those already introduced to explain the flavour structure of the charged fermion sector. Moreover, there is no need to impose small couplings nor assuming that the scales $v_{12}$ or $v_{23}$ are very heavy in order to achieve small neutrino masses (meaning that such scales can be in the multi-TeV region), as is commonly assumed in the literature.
@@ 4.1 Gauge boson couplings and complete fermion mixing
$g_{12}=\frac{g_{1}g_{2}}{\sqrt{g_{1}^{2}+g_{2}^{2}}},\qquad\qquad g_{Y}=\frac{g_{12}g_{3}}{\sqrt{g_{12}^{2}+g_{3}^{2}}}\,.$ (59)
@@ 4.2 Bounds from FCNCs
In contrast to $Z^{\prime}_{12}$, the $Z^{\prime}_{23}$ boson is mostly protected from the most dangerous FCNCs by an approximate $U(2)^{5}$ flavour symmetry. The exception to this appears in Model 1 only, where we find that $Z^{\prime}_{23}$ mediates a 1-loop contribution to $\mu\rightarrow e\gamma$ with $E_{23}$ running in the loop, as shown in Fig. 5 (right), that appears due to the mixing of the $E_{23}$ VL fermion [^8] This is a consequence of non-minimal breaking of the approximate $U(2)^{5}$ flavour symmetry by the messengers $F_{23}$ in Model 1. This is avoided in the (less minimal) model of FernandezNavarro:2023hrf , where the leading breaking of $U(2)^{5}$ is done minimally, SM-like via the VL fermion $Q_{3}\sim(\mathbf{3,2})_{(0,0,\frac{1}{6})}$ and heavy Higgs doublets $H_{2}^{u,d}\sim(\mathbf{1,2})_{(0,\frac{1}{2},0)}$. Alternatively, in Model 2 the $U(2)^{5}$ symmetry is broken minimally in the down and charged lepton sectors, and non-minimally in the up sector where experimental bounds are much weaker. For recent revisits of $U(2)$ symmetries we refer to Greljo:2023bix ; Antusch:2023shi ; Greljo:2024zrj . with $\mu$ and $e$. This contribution to $\mu\rightarrow e\gamma$ is chirally enhanced by a factor $M_{E_{23}}/m_{\mu}\sim 10^{7}$ but also suppressed by small mixing angles $s_{26}^{e_{R}}\sim\mathcal{O}(\langle\phi_{\ell 23}\rangle/M_{E_{23}}\sim\lambda^{3})$ and $s_{16}^{e_{L}}\sim\mathcal{O}(s_{12}^{e_{L}}\langle H_{3}^{d}\rangle/M_{E_{23}}\sim 10^{-7})$. As shown in Table 6 and Fig. 6(a), this contribution leads to a bound in the few TeV region, stronger than the bounds obtained from $3\rightarrow 2$ flavour transitions such as $B_{s}$ meson mixing or $B_{s}\rightarrow\mu\mu$. In Fig. 6(a) we also depict in green the region of parameter space that could explain the anomaly in $b\rightarrow s\nu\nu$ transitions recently observed by Belle II Belle-II:2023esi , although this region is excluded by electroweak precision and by the leading flavour-violating observables. Similarly, we expect significant enhancements of $b\rightarrow s\mu\mu$ or $b\rightarrow s\tau\tau$ (see e.g. Capdevila:2017iqn ; Alguero:2023jeh ; Bordone:2023ybl ) to be in conflict with electroweak precision and with the leading flavour observables.
@@
This bound is well compatible with the model-independent phenomenology of $Z^{\prime}_{23}$, which is dominated by the production of dilepton tails at the LHC and the modification of electroweak precision observables via mixing [^9] As shown in FernandezNavarro:2023rhv , in the tri-hypercharge theory there exists $Z-Z^{\prime}_{23}$ mass mixing even when kinetic mixing among the three gauge $U(1)$ factors is negligible, as it happens when they originate from a semi-simple embedding like in FernandezNavarro:2023hrf . For this reason, in our analysis we neglect kinetic mixing, which is much smaller than the leading mass mixing effect that we include. with the SM $Z$ boson FernandezNavarro:2023rhv ; Davighi:2023evx . In particular, assuming all Yukawa couplings of the theory to be of $\mathcal{O}(1)$, and assuming $g_{3}(10\,\mathrm{TeV})\simeq 0.6$ as suggested by a GUT origin FernandezNavarro:2023hrf , we obtain the same bound $M_{Z^{\prime}_{23}}>5\,\mathrm{TeV}$ from both $\mu\rightarrow e\gamma$ and from dilepton searches at the LHC. Therefore, to be conservative we set $v_{23}\sim\mathcal{O}(10\,\mathrm{TeV})$ and, thanks to the predictive relation in Eq. (25), this fixes the whole spectrum of both models as in Fig. 2, and in particular the mass of $Z^{\prime}_{12}$ as $M_{Z^{\prime}_{12}}\sim v_{12}\sim\mathcal{O}(10^{3}\,\mathrm{TeV})$, well compatible with all the bounds shown in Table 6.
@@ Appendix C EFT formalism for FCNCs
Now we apply a Fierz transformation to the LR operators in order to match to the basis of Refs. UTfit:2007eik ; Bona:2024bue ; UTfit2023 , from which we take the bounds,
$\mathcal{L}_{\mathrm{EFT}}\supset-\left[C_{1}^{sd}(\bar{s}_{L}\gamma_{\mu}d_{L})^{2}+C_{5}^{sd}(\bar{s}_{L}^{\alpha}d_{R}^{\beta})(\bar{s}_{R}^{\beta}d_{L}^{\alpha})+C_{1}^{cu}(\bar{c}_{L}\gamma_{\mu}u_{L})^{2}+C_{5}^{cu}(\bar{c}_{L}^{\alpha}u_{R}^{\beta})(\bar{c}_{R}^{\beta}u_{L}^{\alpha})\right]\,,$ (85)
obtaining $C_{1}^{sd,cu}=C_{LL}^{sd,cu}$ and $C_{5}^{sd,cu}=-2C_{LR}^{sd,cu}$.